OpenStudy (anonymous):

rove the identiy of cos 2u=2 cos^2u-1

OpenStudy (anonymous):

So we have cos 2u = Cos(U + U) If U = A, U = B Then cos(A+B) = cosAcosB - sinAsinB (known identity) = cos U. cos U - sin U.sinU => cos2U - sin2U => cos2U - (1 - cos2U) (sin2U + cos2U = 1) => 2cos2U - 1 = cos 2U. proved.